Use the data in the following​ table, which lists​ drive-thru order accuracy at popular fast food chains. Assume that orders are randomly selected from those included in the table. ​

Drive-thru Restaurant  /  A  /  B  /  C  /  D
Order Accurate         / 47 / 5 / 49 / 1
Order Not Accurate     / 45 / 7 / 12 / 45

If three different orders are​ selected, find the probability that they are all from restaurant B. Assume that the orders are selected without replacement. Express your answer as a percentage rounded to the nearest hundredth.
Answer #1

Ans:

A B C D Total
Accurate 47 5 49 1 102
Not accurate 45 7 12 45 109
Total 92 12 61 46 211

There are total 5+7=12 orders from restaurant B.

There are total 211 orders.

P(restaurant B)=(12/211)*(11/210)*(10/209)=0.0001 or 0.01%